OpenStudy (anonymous):

determine the poits at which the graph of the function has a horizontal tangent: f(x)=x^2/x^2+1

jimthompson5910 (jim_thompson5910):

what is f ' (x) ?

OpenStudy (anonymous):

not sure...

jimthompson5910 (jim_thompson5910):

you have to use the quotient rule

OpenStudy (anonymous):

would i have to find it with the quotient rule ?

OpenStudy (anonymous):

ooh okay

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

\[2x(x^2+1)-(x^2)(2x+1) / (x^2 +1)^2\] then i cancel out the top and bottom (x^2+1) ?

jimthompson5910 (jim_thompson5910):

you can't do that cancellation

jimthompson5910 (jim_thompson5910):

also, you should have 2x(x^2+1) - (x^2)*2 in the numerator

jimthompson5910 (jim_thompson5910):

that simplifies to 2x*x^2 + 2x*1 - x^2*2 2x^3 + 2x - 2x^3 2x

jimthompson5910 (jim_thompson5910):

after everything is simplified, you are left with \[\Large f^{\prime}(x) = \frac{2x}{(x^2+1)^2}\]

OpenStudy (anonymous):

it is in the numerator :o the denominator is the (x^2+1)^2 ... i typed it wrong ?

jimthompson5910 (jim_thompson5910):

you had 2x+1 when it should be 2x sorry I meant to say 2x and not just 2

OpenStudy (anonymous):

ooh okay i see !

jimthompson5910 (jim_thompson5910):

once you have \[\Large f^{\prime}(x) = \frac{2x}{(x^2+1)^2}\], you replace f ' (x) with 0 then solve for x

OpenStudy (anonymous):

so it'll be one ?

jimthompson5910 (jim_thompson5910):

nope

jimthompson5910 (jim_thompson5910):

\[\Large f^{\prime}(x) = \frac{2x}{(x^2+1)^2}\] \[\Large 0 = \frac{2x}{(x^2+1)^2}\] solve for x

OpenStudy (anonymous):

0 ?

jimthompson5910 (jim_thompson5910):

yes x = 0 is a solution, so the horizontal tangent is at (0,0)

jimthompson5910 (jim_thompson5910):

you find y = 0 by plugging x = 0 back into f(x)

OpenStudy (anonymous):

i struggle so much with ap calculus ;( do you have any advice that maybe i can take ??

jimthompson5910 (jim_thompson5910):

well the key when you struggle with any material is to figure out what you know well already

jimthompson5910 (jim_thompson5910):

then try to connect the new stuff to the old stuff you know

jimthompson5910 (jim_thompson5910):

from my experience, calculus has a lot of algebra in it (example: simplifying complicated expressions). So if you struggle with topics like that, then I recommend you practice more with it. As for the complicated derivative formulas, you can come up with clever ways to memorize them. Like with the quotient rule, you can use this trick http://www.howtoace.com/HTACFiles/node21.html or I like to think of it as a slightly modified form of the product rule (numerator' - denominator'). However works for you is fine

OpenStudy (anonymous):

okay so whenever i am finding the tangent of a line i need to use the quotient rule ?

jimthompson5910 (jim_thompson5910):

no, you use the quotient rule when deriving functions of the form f(x) = g(x)/h(x)

jimthompson5910 (jim_thompson5910):

you're given f(x)=x^2/x^2+1 so g(x) = x^2 and h(x) = x^2 + 1

OpenStudy (anonymous):

ooh okay okay. and you always need to find the derivative in order to find the tangent ?

jimthompson5910 (jim_thompson5910):

when they say "find the slope of the tangent line", you have to find f ' (x) aka you have to differentiate or derive f(x) to get f ' (x)

jimthompson5910 (jim_thompson5910):

the derivative f ' (x) is used to find the slope of the tangent line

jimthompson5910 (jim_thompson5910):

yes that is correct

OpenStudy (anonymous):

okay so whenever i need to find the tangent of a line i first have to find the derivative by using the quotient rule correct ?

jimthompson5910 (jim_thompson5910):

"okay so whenever i need to find the tangent of a line i first have to find the derivative" correct "... by using the quotient rule" only if you're dealing with a fraction of two functions

jimthompson5910 (jim_thompson5910):

if you had f(x) = x^2 you just derive using the power rule (no need to use the quotient rule)

OpenStudy (anonymous):

and if it isnt i use the power rule

OpenStudy (anonymous):

okay got it

jimthompson5910 (jim_thompson5910):

there are other rules as well. not just these two rules

jimthompson5910 (jim_thompson5910):

the power rule is for power functions a power function is in the form y = a*x^b 'a' and 'b' are constants

OpenStudy (anonymous):

so a and b would become 0 ?

jimthompson5910 (jim_thompson5910):

well in the case of y = x^2, which is really y = 1*x^2, a = 1 and b = 2

jimthompson5910 (jim_thompson5910):